# NeurDB: On the Design and Implementation of an AI-powered Autonomous Database

### v1

@@ 4.2. Fast-adaptive Learned Components
In this section, we introduce two key learned system components, including learned concurrency control and learned query optimizer, which can achieve fast adaptation to data and workload drift. To achieve this, we fundamentally rely on the underlying AI ecosystem. We non-intrusively monitor the system conditions such as transaction throughput and data distributions, which can be used to detect data and workload drift in real time. Based on these metrics, NeurDB automatically triggers model fine-tuning to adapt to continuously evolving data and workloads, and continually generates valid input for model pre-training, allowing the model for learned system components to gain global knowledge of most drifts.
Learned Concurrency Control
We design an efficient learned concurrency control algorithm that can continually adapt to changing workloads, as shown in Figure 4. Given a transaction $T$ with multiple operations, our algorithm incorporates a decision-making model $\mathcal{{F}}$ to assign each operation $op$ the optimal concurrency control action $\delta$ based on the current system condition represented as $x$, i.e., $\delta=\mathcal{{F}}(x)$. Unlike the state-of-the-art approach Wang et al. (2021) that simply adjusts actions based on predefined transaction or operation patterns (e.g., transaction type), our approach learns the optimal action based on the contention state, which includes both conflict information (such as dependency) of transactions and contextual information (such as the transaction length). For instance, when a write is performed on high-contention data records, we may immediately abort the transaction to avoid unnecessary costs, as the transaction is likely to be aborted eventually. In contrast, we may execute a read on low-contention records with optimistic concurrency control without locking to avoid extra conflict detection overhead, and a long transaction that is supposed to run for a long time should have high priority. By using a modelling paradigm based on the contention state, our model is more likely to generalize to drifting workloads with varying levels of contention. However, since transactions can be completed in milliseconds, the model must be efficient so as not to become a bottleneck. To achieve this, we first develop a fast encoding technique to significantly reduce the dimension of contention state representation, and then compress the model with a flattened layer to improve inference efficiency.
@@
Learned Query Optimizer
We propose a learned query optimizer that can efficiently adapt to data and workload drift, as shown in Figure 5. In contrast to existing works Marcus et al. (2021); Zhu et al. (2023) which attempt to produce the best plan for a given query $\mathcal{Q}$ under fixed system conditions (i.e., data distribution and workload), our approach effectively identifies the plan best suited for the current system conditions. To achieve this, we design a dual-module model consisting of an encoder and an analyzer. Specifically, in the encoder, we input the vector generated by a tree transformer, which includes multiple candidate query plans for $\mathcal{{Q}}$, along with system condition representations (including buffer information depicting buffer usage and data statistics representing each attribute’s distribution), into cross-attention layers to generate a unified embedding. The analyzer then uses this embedding as input for a multi-head attention layer followed by a multilayer perceptron (MLP) to deliver the optimal plan. With more knowledge of the mapping between plans and system conditions, the model can better generalize to data and workload drifts. To maximize this knowledge, we generate various synthetic data distributions and workloads using Bayesian optimization, and pre-train the model to handle most drifts effectively. Consequently, the proposed learned query optimizer provides consistent query performance under evolving data and workloads.
@@ 5. Evaluation
In this section, we present our preliminary evaluation results. We first introduce the experimental setup, and then evaluate the in-database AI ecosystem under real-world AI analytics scenarios, and the learned system components with data and workload drifts.
@@ 5.1.1. Benchmarks
We construct two real-world applications that require AI analytics.
E-commerce (E) Workload. We perform click-through rate prediction, a critical task in e-commerce for product recommendations, using the Avazu dataset (Avazu) ava (2024), which consists of ~40.4M records and 22 attributes. We use k-means clustering to create five data clusters namely, $C_{1}$ to $C_{5}$, and by switching from one to another, we simulate the data distribution drift.
Healthcare (H) Workload. We conduct disease progression prediction using the UCI Diabetes dataset (Diabetes) dia (2024). After scaling, the dataset comprises ~5.2M data records and 43 attributes.
We establish a micro-benchmark to evaluate the learned system components. It consists of a transactional benchmark based on YCSB Cooper et al. (2010), which generates synthetic workloads for large-scale Internet applications. Each transaction performs 5 selects and 5 updates on a table with 1 million records. In addition, we construct an OLAP benchmark based on the STATS dataset sta (2024), which consists of 8 tables from the Stats Stack Exchange network. We execute inserts/updates/deletes with randomly generated data values to simulate data distribution drift following a recent work Li et al. (2023b).
@@ 5.1.2. Implementation and Default Configuration
We have released the first version of NeurDB in neu (2024), in which we integrate our proposed in-database AI ecosystem and learned components into the codebase of PostgreSQL v16.3. We will gradually introduce new modules such as AI-powered resource scheduling, etc., and replace existing components where necessary. We implement a baseline system called PostgreSQL+P, which loads data from PostgreSQL in batches, and utilizes an AI runtime built with PyTorch to support AI analytics. By default, we employ ARM-Net Cai et al. (2021), an adaptive relation modelling network tailored for structured data, as the basic analytics model for both PostgreSQL+P and NeurDB. In our experiments, we inherit the default settings of PostgreSQL unless otherwise specified. We set the default window size of the streaming data loader to 80 data batches. Each batch contains 4096 data records (samples).
@@ 5.2. In-database AI Analytics
5.3. Learned System Components
We now investigate the performance of our proposed learned system components.
@@ 5.3. Learned System Components
To evaluate the proposed concurrency control algorithm, we compare NeurDB with PostgreSQL using the micro-benchmark with varying thread counts. The results, shown in Figure 7(a), demonstrate that NeurDB achieves up to 1.44$\times$ higher transaction throughput than PostgreSQL. We attribute this performance gain to the proposed algorithm’s ability to schedule transactions more effectively than serializable snapshot-isolation Ports and Grittner (2012), a static concurrency control algorithm employed in PostgreSQL. We further evaluate the adaptability of our proposed algorithm against Polyjuice Wang et al. (2021), the state-of-the-art learned algorithm. Due to the limitation of constraining the transaction execution workflow, implementing Polyjuice on the codebase of NeurDB would be cumbersome. We therefore opt to implement our algorithm, named NeurDB(CC), into the Polyjuice codebase to facilitate a fair comparison. We set up a drift workload based on TPCC Council (2007) by varying the number of warehouses and threads. As shown in Figure 7(b), NeurDB(CC) adapts quickly to workload drifts and outperforms Polyjuice by up to 2.05$\times$. The superior performance of NeurDB(CC) mainly stems from its design that encapsulates a fast yet accurate model to find the best concurrency control action, while facilitating the fine-tuning process with the two-phase adaptation algorithm.
@@
As can be observed in Figure 8, NeurDB achieves up to 20.32% lower average latency of all evaluated queries, which demonstrates its effective adaptability to both data and workload drifts. Due to the proposed dual-module model, NeurDB is able to effectively capture system conditions and use them to select an efficient query plan.
@@ 6. Related Work
Our work relates to a broad spectrum of efforts on the fusion of AI and databases. The initial concepts Brodie (1988); Graefe and DeWitt (1987) can be traced back to the 1980s. At that time, the fusion was far from reality due to limited development in both realms. With advancements in both fields over the years, numerous attempts by academia  Kraska et al. (2019); Pavlo et al. (2017) and industry Karanasos et al. (2020); Li et al. (2021) have been made. However, many problems remain open, and it is important to consider database dynamicity in the fusion of AI with databases. Further, the system architecture may require a re-design to support AI-powered operations. NeurDB has been designed along this direction by seamlessly integrating AI into the database architecture, enabling continuous adaptation to data and workload drift.
@@ 7. Conclusions
This paper presents NeurDB, a novel AI-powered autonomous DBMS that is adaptable to data and workload drift. NeurDB, with its fast-adaptive learned system components and in-database AI ecosystem, facilitates efficient and effective in-database AI analytics. Empirical evaluations demonstrate the superiority of NeurDB, highlighting its potential to realize a seamless fusion of AI and databases.

### v2

@@ 4.2. Fast-adaptive Learned Components
In this section, we introduce two key learned database components, learned concurrency control and learned query optimizer, which can achieve fast adaptation to data and workload drift. To achieve this, we fundamentally rely on the underlying AI ecosystem. We non-intrusively monitor the system conditions such as transaction throughput and data distributions, which can be used to detect data and workload drift in real time. Based on these metrics, NeurDB automatically triggers model fine-tuning to adapt to continuously evolving data and workloads, and continually generates valid input for model pre-training, allowing the model for learned database components to gain global knowledge of most drift.
Learned Concurrency Control
We design an efficient learned concurrency control algorithm that can continually adapt to changing workloads, as shown in Figure 4. Given a transaction $T$ with multiple operations, our algorithm incorporates a decision-making model $\mathcal{{F}}$ to assign each operation $op$ the optimal concurrency control action $\delta$ based on the current system condition represented as $x$, i.e., $\delta=\mathcal{{F}}(x)$. Unlike state-of-the-art approach Wang et al. (2021) that simply adjusts actions based on predefined transaction or operation patterns (e.g., transaction type), our approach learns the optimal action based on the contention state, which includes both conflict information (such as dependency) of transactions and contextual information (such as the transaction length). For instance, when a write is performed on high-contention data records, we may immediately abort the transaction to avoid unnecessary costs, as the transaction is likely to be aborted eventually. In contrast, we may execute a read on low-contention records with optimistic concurrency control without locking to avoid extra conflict detection overhead, and a long transaction that is supposed to run for a long time should have high priority. By using a modeling paradigm based on the contention state, our model is more likely to generalize to drifting workloads with varying levels of contention. However, since transactions can be completed in milliseconds, the model must be efficient so as not to become a bottleneck. To achieve this, we first develop a fast encoding technique to significantly reduce the dimension of contention state representation, and then compress the model with a flattened layer to improve inference efficiency.
@@
Learned Query Optimizer
We propose a learned query optimizer that can efficiently adapt to data and workload drift, as shown in Figure 5. In contrast to existing works Marcus et al. (2021); Zhu et al. (2023) that attempt to produce the best plan for a given query $\mathcal{Q}$ under fixed system conditions (i.e., data distribution and workload), our approach effectively identifies the plan best suited for the current system conditions. To achieve this, we design a dual-module model consisting of an encoder and an analyzer. Specifically, in the encoder, we input the vector generated by a tree transformer, which includes multiple candidate query plans for $\mathcal{{Q}}$, along with system condition representations (including buffer information depicting buffer usage and data statistics representing each attribute’s distribution), into cross-attention layers to generate a unified embedding. The analyzer then uses this embedding as input for a multi-head attention layer followed by a multilayer perceptron (MLP) to deliver the optimal plan. With more knowledge of the mapping between plans and system conditions, the model can better generalize to data and workload drift. To maximize this knowledge, we generate various synthetic data distributions and workloads using Bayesian optimization, and pre-train the model to handle most drift effectively. Consequently, the proposed learned query optimizer provides consistent query performance under evolving data and workloads.
@@
Discussion
The main design principle we follow to develop computational models and algorithms in NeurDB is the filter-and-refine principle (FRP) Ooi et al. (2024, 2002); Filter and Refine (2024). FRP employs a two-stage strategy: a filtering stage quickly eliminates less promising or irrelevant objects (e.g., data, strategies, etc.) from a large set, followed by a more resource-intensive refinement stage on the filtered subset. This strategy enables the efficient execution of computationally demanding tasks while yielding near-optimal results. We therefore leverage FRP in both in-database AI analytics and autonomous DBMS optimization for enhanced adaptability. For example, we integrate FRP in the RL framework of learned concurrency control and incorporate it in the multi-head attention layers of the learned query optimizer. However, existing model structures, typically characterized by fixed topologies and layer configurations, constrain the full potential of FRP. We are exploring a flexible model structure that can be dynamically adjusted with the guidance of FRP, enabling seamless adaptation to data and workload drift.
@@ 5. Evaluation
In this section, we present our preliminary evaluation results. We first introduce the experimental setup, and then evaluate the in-database AI ecosystem under real-world AI analytics scenarios, and the learned database components with data and workload drifts.
@@ 5.1.1. Benchmarks
We construct two real-world applications that require AI analytics. The AI analytics queries used in our experiments are listed in Table 1.
  • E-commerce (E) Workload performs click-through rate prediction, a critical task in e-commerce for product recommendations, using the Avazu dataset (Avazu) Dataset (2024a), which consists of ~40.4M records and 22 attributes. We use k-means clustering to create five data clusters namely, $C_{1}$ to $C_{5}$, and by switching from one to another, we simulate the data distribution drift.
  • Healthcare (H) Workload conducts disease progression prediction using the UCI Diabetes dataset (Diabetes) Dataset (2024c). After scaling, the dataset comprises ~5.2M data records and 43 attributes.
We establish a micro-benchmark to evaluate the learned database components. It consists of a transactional benchmark based on YCSB Cooper et al. (2010), which generates synthetic workloads for large-scale Internet applications. Each transaction performs 5 selects and 5 updates on a table with 1 million records. In addition, we construct an OLAP benchmark based on the STATS dataset Dataset (2024b), which consists of 8 tables from the Stats Stack Exchange network. We execute inserts/updates/deletes with randomly generated data values to simulate data distribution drift following a recent work Li et al. (2023a).
Table 1. Queries for AI Analytics Evaluations
Workload Statement
E-Commerce (E) PREDICT VALUE OF click_rate FROM avazu TRAIN ON *
Healthcare (H) PREDICT CLASS OF outcome FROM diabetes TRAIN ON *
@@ 5.1.2. Implementation and Default Configuration
We have released the first version of NeurDB NeurDB (2024), in which we integrate our proposed in-database AI ecosystem and learned components into the codebase of PostgreSQL v16.3. We will gradually introduce new modules such as AI-powered resource scheduling, etc., and replace existing components where necessary. We implement a baseline system called PostgreSQL+P, which loads data from PostgreSQL in batches, and utilizes an AI runtime built with PyTorch to support AI analytics. By default, we employ ARM-Net Cai et al. (2021), an adaptive relation modeling network tailored for structured data, as the basic analytics model for both PostgreSQL+P and NeurDB. In our experiments, we inherit the default settings of PostgreSQL unless otherwise specified. We set the default window size of the streaming data loader to 80 data batches. Each batch contains 4096 data records (samples).
@@ 5.2. In-database AI Analytics
5.3. Learned Database Components
We now investigate the performance of our proposed learned database components.
@@ 5.3. Learned Database Components
To evaluate the proposed concurrency control algorithm, we compare NeurDB with PostgreSQL using the micro-benchmark with varying thread counts. The results, shown in Figure 7(a), demonstrate that NeurDB achieves up to 1.44$\times$ higher transaction throughput than PostgreSQL. We attribute this performance gain to the proposed algorithm’s ability to schedule transactions more effectively than serializable snapshot isolation Ports and Grittner (2012), a static concurrency control algorithm employed in PostgreSQL. We further evaluate the adaptability of our proposed algorithm against Polyjuice Wang et al. (2021), state-of-the-art learned algorithm. Due to the limitation of constraining the transaction execution workflow, implementing Polyjuice on the codebase of NeurDB would be cumbersome. We therefore opt to implement our algorithm, named NeurDB(CC), into the Polyjuice codebase to facilitate a fair comparison. We set up a drift workload based on TPCC TPCC (2024) by varying the number of warehouses and threads. As shown in Figure 7(b), NeurDB(CC) adapts quickly to workload drift and outperforms Polyjuice by up to 2.05$\times$. The superior performance of NeurDB(CC) mainly stems from its design that encapsulates a fast yet accurate model to find the best concurrency control action, while facilitating the fine-tuning process with the two-phase adaptation algorithm.
@@
As can be observed in Figure 8, NeurDB achieves up to 20.32% lower average latency of all evaluated queries, which demonstrates its effective adaptability to both data and workload drift. Due to the proposed dual-module model, NeurDB is able to effectively capture system conditions and use them to select an efficient query plan.
@@ 6. Related Work
Our work relates to a broad spectrum of efforts on the fusion of AI and databases. The initial concepts Brodie (1988); Graefe and DeWitt (1987) can be traced back to the 1980s. At that time, the fusion was far from reality due to limited development in both realms. With advancements in AI and database fields over the years, numerous attempts by academia Kraska et al. (2019); Pavlo et al. (2017) and industry Karanasos et al. (2020); Li et al. (2021) have been made to advance in-database AI analytics and autonomous DBMS optimization. However, many problems remain open, and it is important to consider data and workload drift in both directions. Here we shall briefly discuss recent advancements in AI and DBMS integration, with a particular focus on their approaches to handling data and workload drift.
In-database AI Analytics
Recent research increasingly focuses on optimizing the performance of in-database AI analytics. For example, some studies Salazar-Díaz et al. (2024); Park et al. (2022); Karanasos et al. (2020); Jasny et al. (2020) extend the existing query and transaction execution framework to improve the efficiency of analytics tasks. Further, in-database model management systems Su et al. (2024); Vartak (2017); Zhou et al. (2022) are proposed to minimize model storage costs while efficiently serving models for analytics tasks. These optimizations can be broadly adopted by full-fledged DBMSs supporting AI analytics, such as nsDB Yuan et al. (2024), GaussML Li et al. (2024), and NeurDB. However, apart from the mentioned approaches, NeurDB establishes an in-database AI ecosystem, specifically designed to enhance the efficiency and effectiveness of AI analytics under continuous data and workload drift.
Autonomous DBMS Optimization
Traditional AI-driven DBMS optimization in areas such as knob tuning Pavlo et al. (2017), resource management Li et al. (2023b, 2021), and learned system components Kraska et al. (2019); Marcus et al. (2021); Zhu et al. (2023); Wang et al. (2021); Kraska et al. (2018) typically relies on complete retraining to handle data and workload drift. Recently, the development of drift-handling mechanisms become popular, particularly for learned query optimizers and learned indexes. For example, some works introduce methods to automatically detect drift and trigger model updates Wu and Ives (2024); Li et al. (2022); Kurmanji and Triantafillou (2023), enabling learned query optimizers to withstand data and workload drift. Further, updatable learned indexes Zhang et al. (2024); Ding et al. (2020); Wu et al. (2021) are enhanced with the adaptability to continuous transactional updates. Unlike existing works focusing on certain components, NeurDB introduces a unified architecture based on the proposed in-database AI ecosystem to facilitate fast adaptation across all its key components.
We are extending NeurDB to be an intelligent cloud-native HTAP database with built-in support for advanced AI analytics. In particular, we disaggregate the system architecture into separate compute and storage layers Pang and Wang (2024); Cao et al. (2021), enabling dedicated compute nodes to handle OLTP, OLAP, and AI analytics independently. Based on this foundation, we are devising AI-powered resource scheduling to enable on-demand resource allocation and component assembling, thus enhancing system scalability and adaptability.
@@ 7. Conclusions
This paper presents NeurDB, a novel AI-powered autonomous DBMS that is adaptable to data and workload drift. NeurDB, with its fast-adaptive learned database components and in-database AI ecosystem, facilitates efficient and effective in-database AI analytics and autonomous system optimization. Empirical evaluations demonstrate the superiority of NeurDB, highlighting its potential to realize a seamless fusion of AI and databases.
{acks}
We would like to thank Jiaqi Zhu for her help and comments. The work of NUS researchers is partially supported by the Lee Foundation in terms of Beng Chin Ooi’s Lee Kong Chian Centennial Professorship fund and NUS Faculty Development Fund. Gang Chen’s work is supported by National Key Research and Development Program of China (2022YFB2703100). Yanyan Shen’s work is supported by National Key Research and Development Program of China (2022YFE0200500). Yuncheng Wu’s work is supported by National Key Research and Development Program of China (2023YFB4503600). Meihui Zhang’s work is supported by National Natural Science Foundation of China (62072033).